Given that A(2,3,-1) and B(1,2,4) are points in space. Find the unit vector along the direction of AB and the one along BA. Also find the position vector of the point C which is on the mid-point of AB

1
Expert's answer
2022-04-11T12:53:27-0400

The vector "\overrightarrow{AB}" has the form: "\overrightarrow{AB}=(-1,-1,5)". The length of the vector is: "|\overrightarrow{AB}|=\sqrt{27}". Thus, the unit vector along the direction of "AB" has the form: "(-\frac{1}{\sqrt{27}},-\frac{1}{\sqrt{27}},\frac{5}{\sqrt{27}})". The opposite unit vector has the form: "(\frac{1}{\sqrt{27}},\frac{1}{\sqrt{27}},-\frac{5}{\sqrt{27}})". Find coordinates of the point "C:" "C=(2-\frac12,3-\frac12,-1+\frac52)=(1.5,2.5,1.5)". It is not mentioned, what is the origin of the vector. Consider two options: 1. "O=(0,0,0)" is the origin, Then, "\overrightarrow{OC}=(1.5,2.5,1.5)". It is the position vector. 2. Suppose that "A=(2,3,-1)" is the origin. Then, "\overrightarrow{AC}=(-0.5,-0.5,2.5)".